### v1

# AdTEC: A Unified Benchmark for Evaluating Text Quality in Search Engine Advertising

Abstract
With the increase in the more fluent ad texts automatically created by natural language generation technology, it is in the high demand to verify the quality of these creatives in a real-world setting. We propose AdTEC, the first public benchmark to evaluate ad texts in multiple aspects from the perspective of practical advertising operations. Our contributions are: (i) Defining five tasks for evaluating the quality of ad texts and building a dataset based on the actual operational experience of advertising agencies, which is typically kept in-house. (ii) Validating the performance of existing pre-trained language models (PLMs) and human evaluators on the dataset. (iii) Analyzing the characteristics and providing challenges of the benchmark. The results show that while PLMs have already reached the practical usage level in several tasks, human still outperforms in certain domains, implying that there is significant room for improvement in such area.
1 Introduction
Online advertising, especially sponsored search advertising, is a dominant sector for vendors to promote their products, and the market size is estimated to grow by billions of dollars over the next few years [1]. To meet the increasing demands of advertising operations (AdOps), such as creating ad texts from product information (Step 2 in Figure 1), the remarkable success of natural language generation (NLG) by pre-trained language models (PLMs) [2, 3, 4, 5] has given a boost to applications in practice [6, 7, 8], making advertising a huge industrial use case for NLP.
In this paper, we focus on evaluating the quality of ad texts generated by such models. We refer to this process as ad text evaluation, as depicted in Step 3 of Figure 1. Ad text evaluation is crucial because low-quality ad texts can disadvantage advertisers through a lack of fluency, inappropriate appeals, and misleading representations. Because it is expensive and non-scalable to ask humans for verifying the quality of each text in high-volume domains such as sponsored search advertising, it is in the high demand for developing automatic quality estimators for ad texts. The quality has multiple dimensions, such as appropriate wording, effective appeals, consistency between ad text and product information, and high predicted performance. Although these dimensions should be included for evaluating automatic quality estimators, there is no such benchmark. Consequently, the bottleneck lies in verifying the quality of ad texts despite the ability of generating numerous creatives automatically, which hinders the scalability of delivery volume. Thus, we aim to construct a benchmark for evaluating the quality of ad texts.
The primary challenge in constructing an ad text evaluation benchmark is the absence of a clear definition for the tasks [9, 10]. The lack of domain knowledge in AdOps complicates the understanding of high-quality ad text standards and the accurate definition of tasks. However, the AdOps workflow is complex, relies on various platforms and formats, and encompasses multiple methodologies and metrics. In addition, only a few companies possess the expertise to operate online advertising at scale. Owing to legal and contractual obligations, advertising workflows and data are predominantly managed in-house, leading to a lack of publicly available datasets. This scarcity makes it challenging to systematically reproduce and validate diverse methodologies in academia. Consequently, research remains less active in the advertising field, potentially overlooking issues and delaying the application and development of cutting-edge technology.
In response to these challenges, we propose AdTEC (Ad Text Evaluation Benchmark by CyberAgent), the first publicly available benchmark that defines and unifies tasks based on generalized AdOps workflows. Our major contributions are as follows:
- The First Public Dataset on Ad Text Evaluation.
We carefully designed five tasks that closely match real-world advertising workflows based on practical operational experiences of advertising agencies. Based on these tasks, we constructed a high-quality Japanese dataset from the operational data and released it publicly, even though such data is typically being kept in-house and difficult to obtain. These datasets are available at https://github.com/CyberAgentAILab/AdTEC.
- Benchmark Experiments.
We validated the performance of existing PLMs, such as BERT, RoBERTa, and large language models (LLMs), as well as human evaluators on our proposed benchmark.
- Dataset Analysis.
We analyzed the characteristics and identified potential issues of the dataset through experiments, demonstrating that our benchmark is challenging, and highlighting potential areas for improvement and future research.
2 Understanding Sponsored Search Advertising and its Workflow
Sponsored search advertising displays titles and descriptions of ads that are relevant to keywords, which users enter into search engines, appearing as part of the search results. As illustrates in Figure 2, when a user clicks ad’s URL, they are directed to a web page known as a landing page (LP). LP contains texts and images related to the advertiser’s products, allowing the user to take actions, such as making a purchase.
Operating such ads requires high expertise, owing to the variety of delivery platforms, properties and formats. To gain insights, we interviewed two types of experts familiar with AdOps; those overseeing AdOps departments at ad agencies and those directly involved in on-site operations. Through these interviews, we explored the intricacies of AdOps, generalized the workflow, and delineated six steps, as illustrated in Figure 1. In this work, we focus mainly on Step 3 of quality evaluation.
3 Building AdTEC
Our goal is to develop a benchmark that captures multi-dimensional aspects of quality that are relevant to practical scenarios and carefully curated to provide value for both real-world applications and various research purposes.
3.1 Task Design
We followed the workflow outlined in §2, and asked the same experts from the interview to identify crucial parts within Step 3 in Figure 1. Derived from their insights, we generalized and defined tasks that adhere to the principle of evaluating ad texts directly or indirectly. Direct evaluation tasks are used to test texts with strict criteria, such as a binary pass/fail outcome or a numerical score to quantify the text quality. These tasks will serve as a checklist to ensure that minimum delivery standards are met. Indirect evaluation tasks are used to assist human evaluators in reviewing or refining texts or serve as a bridge to connect the downstream tasks. Based on the above principles, we designed five tasks: three direct evaluation tasks (Ad Acceptability, Ad Consistency, and Ad Performance Estimation) and two indirect evaluation tasks (A3 Recognition and Ad Similarity).111Note that all examples in this paper are translated from Japanese into English for the presentation brevity.
| Ad text | LP text | Ad Accept. / Consist. |
| Condominium Sales / Free Assessment Now | Let XX assess your single-family home! | acceptable / inconsistent |
| Engineer’s career / Engineer’s Job | Find jobs at website XX! | unacceptable / consistent |
Ad Acceptability . As most ad delivery platforms impose text length restrictions, minor grammatical errors are tolerated to enhance the readability and engage readers within limited space. However, excessive compression can mislead readers, and such poor-quality ads should be detected before delivery to avoid negative impacts on the advertiser. To assess this, we defined the task of Ad Acceptability, which predicts the acceptance of overall quality with binary labels: acceptable/unacceptable. Unacceptable ad phenomena include “collapsing symbols,” “unnatural repetition,” and “incomprehensible meaning” based on expert feedback. This differs from the general concept of linguistic acceptability, which checks for grammatical correctness, such as CoLA [11]. Examples of Ad Acceptability are shown in Table 1. The ad text “Engineer’s career / Engineer’s Jobs” is unacceptable because the meaning is duplicated.
Ad Consistency . Verifying consistency between the ad text and LP content is crucial. If a feature or price mentioned in the ad text is not referenced in the corresponding LP, it may violate the Law for the Prevention of Unjustified Extra or Unexpected Benefit and Misleading Representation, resulting in damages to the advertiser. However, these inconsistencies are hard to detect as some factual expressions do not appear in LPs. For example, the term “official,” which is often used in ad text but rarely appears in LP content. To assess this, we defined the Ad Consistency task, which predicts the consistency between LP content and ad text with binary labels: consistent/inconsistent. Examples of Ad Consistency are shown in Table 1. The first line is labeled inconsistent because the LP refers to a “single-family home,” while the ad text mentions a “condominium.”
Ad Performance Estimation . The most straightforward way to measure ad quality is to publish them online and let end customers evaluate them. However, delivering all ads without alterations is impractical, as low-quality ads can negatively impact advertisers. Therefore, prior studies investigated offline methods to measure ad text quality by simulating customer behavior, such as click-through rate (CTR), based on past delivery history. These methods are currently standard practices in many organizations. Inspired by their works [12, 13, 14], we adopted Ad Performance Estimation task to estimate a quality score in the range of [0, 100] from ad texts, keywords, and industry, as shown in Table 2. The score simulates customer behavior based on past delivery history, and is non-linearly transformed to maintain the original label distribution for contractual reasons.
A3 Recognition . One of the most crucial factors in advertising is the aspect of advertising appeals (A3). At its core, advertising aims to connect advertisers with readers, and A3 serves as a bridge between them. For example, an ad emphasizing low cost may resonate with price-conscious readers, while one focusing on high performance may not. Thus, recognizing appealing expressions in advertising and using appropriate A3 can enhance downstream tasks, such as CTR prediction [9]. In the A3 Recognition task, we follow a previous study [9], which predicts all relevant A3 labels in a given ad text. Figure 3 shows an example of ad texts and the corresponding A3s. All labels and distributions can be found in Appendix Table 13.
Ad Similarity . Repeatedly showing same ads to readers leads to ad fatigue [15], where readers become bored and ad performance declines. Therefore, it is essential to avoid displaying the same ads for extended periods, and regularly replace them with different ones. However, the transition from old to new ads must be carefully managed because we need to maintain the product and its appeal while updating the wording or representations, or we risk disengaging customers who were attracted to the previous ads. Thus, measuring the similarity particularly focusing on this situation is crucial, which enables us to determine whether to replace the ad based on a quantified score.
Building on the aforementioned motivation, we defined the Ad Similarity task, which predicts the similarity score for an ad text pair on a scale of [1, 5]. The lower the values, the less similar the pair, and vice versa. Examples are shown in Table 3. The first example pair illustrates high similarity, where the both the product of “Suppon Black Vinegar” and the A3 of luxury are identical. Conversely, the second example pair differ in A3 with budget and discount, resulting in relatively low similarity.
| Field Example value Title 1 [No.1] Card loan comparison site Title 2 A must-see for those who in a hurry! Title 3 Instant Loan Secure Card Loan Desc. 1 The best place to get a card loan without telling anyone. You only need a driver’s license to apply Desc. 2 Convenient to use ATMs at convenience stores. Convenient and quick loans are available if you apply before 10:00 p.m. Keyword card loan Industry finance Score 82.3 Table 2: Examples of Ad Performance Estimation task. Desc. represent Description. Score is the label and others are inputs. |
3.2 Dataset Construction
Data Collection. For the Ad Acceptability and Ad Consistency task, data was collected from the ad creation phase of the actual AdOps workflow, including both human creators and NLG models’ outputs. For the Ad Performance Estimation and Ad Similarity task, we used Japanese sponsored search ads delivered between 2021 and 2022. For the A3 Recognition task, we used data from Murakami et al. [9].
Data Pre-processing. In the Ad Performance Estimation task, we negotiated with our clients to use a subset of the data approved for public release, as the data includes ad performance metrics sensitive to advertisers. Furthermore, we applied a nonlinear transformation to the raw CTR, scaling them to the range of to preserve their distribution. Additionally, we masked proper nouns (e.g., product names, company names) to prevent identification of advertisers and avoid any potential negative repercussions upon data release. In the Ad Similarity task, creating sentence pairs through random sampling is inefficient because most pairs are not similar. Therefore, we utilized the account structure configured during ad delivery, which includes client, account, campaign, ad group, and keyword information, as depicted in Figure 4. We created pseudo-similar pairs by sampling texts from the same ad group, as preliminary results suggested. To balance the label distribution, we created pseudo-dissimilar pairs by ensuring that the two texts belonged to different clients. Consequently, we sampled pseudo-similar and pseudo-dissimilar pairs at a ratio of 9:1.
Annotation Workflow. All annotators are native Japanese speakers and experts with vast experience in AdOps. The annotation workflow for all tasks, except Ad Performance Estimation and A3 Recognition, is as follows: (1) We first removed duplicate entries and filtered out data in languages other than Japanese. (2) We iteratively revised the annotation guidelines until we achieved a satisfactory level of agreement through pilot annotations on small sampled datasets. (3) Following the guidelines, we conducted a pilot annotation by asking three annotators to annotate the same sampled data used in Step 2. Thereafter, we compared the annotators’ results with our own and resolved any inconsistencies to further refine the guidelines. This cycle was repeated at least twice. (4) After completing Steps 1 to 3, we conducted the main annotation on the full test set using the finalized guidelines. The complete guidelines are provided in the Appendix.
| Task | Train | Dev | Test |
| Ad Acceptability | 13,265 | 970 | 980 |
| Ad Consistency | 10,635 | 945 | 970 |
| Ad Perf. Est. | 125,087 | 965 | 965 |
| A3 Recognition | 1,856 | 465 | 410 |
| Ad Similarity | 4,980 | 623 | 629 |
Data Splitting. Despite efforts at deduplication, similar ad expressions can still be easily found, potentially leading to data leakage with a simple random split. Furthermore, assuming the data is used in industry, it is crucial to generalize effectively without overfitting to specific ad expressions. Therefore, for Ad Acceptability and Ad Consistency, we split the data, considering the ad hierarchical structure in Figure 4, ensuring that clients do not overlap across training, development, and testing. For the Ad Performance Estimation, we used the delivery structure, with the non-overlapping layer as the campaign. For the A3 Recognition, we used the same split from Murakami et al. [9]. For the Ad Similarity, we randomly split the data to maintain label distribution consistency.
4 Benchmark Experiment Settings
In this section, we conducted an experiment to explore the performance on the proposed benchmark.
4.1 Evaluation Metrics
Table 5 provides a brief overview of each task and the corresponding metrics used to measure task performance. Our tasks are categorized into three setups: binary classification, multi-label classification, and regression. For binary classification, we use accuracy and the F1-score to evaluate the binary labels. In multi-label classification, we follow Murakami et al. [9] and use the F1-score with both macro and micro settings. For regression, we use Pearson and Spearman correlation coefficients. These metrics are standard practices in numerous studies.
| Task | Setup: Input Label | Metrics |
| Ad Acceptability | Classification: Ad text acceptable/unacceptable | Accuracy/F1-score |
| Ad Consistency | Classification: (Ad text, LP text) consistent/inconsistent | Accuracy/F1-score |
| Ad Perf. Est. | Regression: (Ad texts, Keyword, Industry) | Pearson/Spearman corr. |
| A3 Recognition | Classification: Ad text multi-labels (See Appendix Table 13) | F1-micro/-macro |
| Ad Similarity | Regression: Ad text pair | Pearson/Spearman corr. |
4.2 Evaluators
We employed two types of evaluators: the PLMs, including both fine-tuned settings with encoder models and zero-/few-shot settings for LLMs, as well as human evaluators. Please note that fine-tuning an LLM is beyond the scope, owing to the cost constrains.
Fine-tuning Setting with Encoder Models.
In the fine-tuning settings, models receive ad text and other values as input and predict labels , where represents the -th token of the input sequence. The hyperparameters used are shown in Appendix Table 12. Ad Acceptability and Ad Consistency involve predicting binary label using an MLP from the encoded vector representation , where is the special token for classification. The inputs of the Ad Acceptability task only use the ad text , while Ad Consistency uses as input, where and represent the LP text and concatenation by a special token , respectively. A3 Recognition involves predicting all possible labels from a given ad text, . We adopted the doc-based architecture described in Murakami et al. [9], where the encoder model was used to obtain the vector representation . This vector was then fed into an MLP layer, producing a label probability distribution , where and is 21. This number corresponds to the number of A3 labels defined by Murakami et al. [9], as shown in Appendix Table 13. Ad Similarity and Ad Performance Estimation are regression tasks that involve predicting a value of range and , respectively. Similar to the text classification task, we concatenated all inputs for each task with the special token , encoded the input into vector representation , and then fed the vector into the MLP layer to predict .
We utilized publicly available encoder models as baselines, namely Tohoku BERT, Waseda RoBERTa, and XLM-RoBERTa [16], commonly employed for Japanese NLP tasks, which differ in pre-tokenizer, tokenization unit, and pre-training dataset. All the aforementioned models are of LARGE size, with detailed information provided in Appendix Tables 8 and 9. We conducted experiments using the BASE size model, however, due to space constraints, the results from BASE size models are provided in Appendix Table 10.
Zero-/Few-shot Setting with LLMs.
We employed CALM27b and ELYZA7bas baselines for open LLMs. CALM27b and ELYZA7b are based on the Llama 2 [17] architecture but differ in training methods and data; CALM27b was trained from scratch, whereas ELYZA7b was continuously trained from the original Llama 2. These models were chosen because of their public availability and strong performance. Additionally, OpenAI’s GPT-3.5 and GPT-4 were included as reference baselines, owing to their exceptional performance, despite not being publicly available. Note that the results from GPT-3.5 are shown in the complete result in Appendix Table 10.
Before conducting the experiment, we sampled 100 instances from the development set to adjust the parameters and prompts. The prompts we used are shown in Appendix Figure 7. Additionally, we determined the number of shots that performed best in the development set: 3-shot for Ad Similarity and Ad Performance Estimation task, 2-shot for A3 Recognition, and zero-shot for the others. We calculated the final score by averaging five runs with different few-shot examples for a single instance.
Human.
We enlisted three human evaluators who are not engaging in AdOps to evaluate all tasks except Ad Performance Estimation. We followed the same procedure for instruction as described in §3.2. Pilot evaluations were carried out twice on randomly sampled 100 instances from the training set for each task before the main run. In the final assessment, a majority vote per instance was conducted for Ad Acceptability and Ad Consistency tasks, while the average scores of evaluator assessments were reported for other tasks.
| Evaluator | Ad Accept. | Ad Consist. | Ad Perf. Est. | A3 Recognition | Ad Similarity |
| Accuracy/F1-score | Accuracy/F1-score | Pearson/Spearman | F1-micro/-macro | Pearson/Spearman | |
| Fine-tuned Encoder Models | |||||
| Tohoku BERT | 0.711/0.688 | 0.767/0.552 | 0.480/0.497 | 0.774/0.694 | 0.773/0.807 |
| Waseda BERT | 0.598/0.637 | 0.755/0.474 | 0.445/0.457 | 0.663/0.517 | 0.740/0.800 |
| XLM-RoBERTa | 0.705/0.690 | 0.758/0.519 | 0.453/0.457 | 0.778/0.677 | 0.878/0.878 |
| Zero-/Few-shot LLMs | |||||
| CALM27b | 0.520/0.115 | 0.381/0.472 | 0.006/0.013 | 0.154/0.042 | 0.036/0.036 |
| ELYZA7b | 0.352/0.520 | 0.628/0.771 | 0.003/0.046 | 0.196/0.044 | 0.015/-0.004 |
| GPT-4 | 0.325/0.433 | 0.583/0.612 | 0.028/0.073 | 0.417/0.113 | 0.776/0.811 |
| Human | 0.732/0.790 | 0.703/0.807 | — | 0.564/0.538 | 0.699/0.765 |
5 Result and Discussion
Table 6 provides an overview of the result. For the complete result including base size encoder models and GPT-3.5, please refer to Appendix Table 10.
Fine-tuned Setting with Encoder Models. XLM-RoBERTa and Tohoku BERT achieved the highest or competitive scores in two or more tasks. In addition, the LARGE model outperformed the BASE model in most tasks, suggesting that the increasing parameter size plays a key role in understanding ad expressions.
Zero-/Few-shot Setting with LLMs. GPT-4 achieved high performance across all three tasks, while ELYZA7b performed the best among the LLMs in Ad Consistency. A substantial difference was observed between the open and OpenAI’s LLMs in the Ad Similarity task. CALM27b and ELYZA7b scored close to 0, indicating no correlation, whereas OpenAI’s models, especially GPT-4, achieved competitive scores of 0.776/0.811. Thus, the open LLMs used in this study struggle with handling the semantic similarity or numerical answers, whereas GPT-4 performs considerably better. In the Ad Performance Estimation task, all LLMs produced uncorrelated responses close to 0, indicating that accurately predicting ad performance remains a challenge for open LLMs.
Fine-tuned Encoder Models vs. Zero-/Few-shot LLMs. Overall, the fine-tuned models outperformed the LLMs. The difference was substantial, ranging from 0.2 to 0.6, especially for the Ad Acceptability, Ad Performance Estimation, and A3 Recognition tasks. This can be attributed to the characteristics of the tasks; Ad Performance Estimation involves predicting numbers in the range , and A3 Recognition requires selecting all suitable labels from more than 20 labels, suggesting that the variety of data features and outputs could not be handled by few-shot alone. Both Ad Acceptability and Ad Consistency are binary classification tasks, yet the performance gap between fine-tuned models and LLMs is more noticeable in the Ad Acceptability task, with a difference as large as 0.2 points, compared to the relatively smaller difference observed in the Ad Consistency task. The task with the smallest difference was Ad Similarity, with a difference of only 0.06 to 0.10.
PLMs vs. Human. Human evaluators outperform models in Ad Acceptability and Ad Consistency. In both tasks, models tends to have high accuracy and a low F1-score, while humans exhibit the opposite trend. Particularly in the Ad Consistency task, where the label distribution is unbalanced, even the best model achieves an F1-score of only 0.55, whereas human performance reaches 0.8. This suggests that humans can make better predictions in both precision and recall in an unbalanced data. On the other hand, fine-tuned models outperform humans on the A3 Recognition and Ad Similarity tasks. In the A3 Recognition task, human evaluators struggled with the diversity of output labels, similar to LLMs. However, the difference between the F1-micro and F1-macro scores is relatively small at 0.03 points for the human result. In contrast, for the fine-tuned PLMs, the difference ranges from 0.08 to 0.20 points. This indicates that human evaluators have a strong ability to generalize and can maintain higher performance even when a label appears infrequently. In the Ad Similarity task, in addition to fine-tuned models, GPT-4 outperforms human evaluators. This is the only task where LLM outperforms humans, suggesting that GPT-4 has a high level of alignment with humans in terms of semantic similarity and numerical understanding.
| Error | GT | H | M | Accept. | Consist. |
| C-IC | T | T | F | 30.1% | 17.2% |
| F | F | T | 0.6% | 5.5% | |
| IC-C | T | F | T | 0.6% | 1.9% |
| F | T | F | 18.0% | 10.9% | |
| IC-IC | T | F | F | 4.9% | 6.2% |
| F | T | T | 3.9% | 6.6% |
Error Analysis on Ad Acceptability and Ad Consistency task. We have conducted a detailed analysis of two tasks, Ad Acceptability and Ad Consistency, where the model has not yet outperformed human evaluators. We analyzed three types of errors: human incorrect and model correct (IC-C), human correct and model incorrect (C-IC), and both incorrect (IC-IC), as shown in Table 7. In both tasks, the model more often predicted False labels than True labels for both IC-C and C-IC errors.. In contrast, humans provided more True labels in both correct and incorrect cases. This suggests that humans exhibit a relatively higher degree of leniency compared to models, which tend to be overly cautious when making decisions in Ad Acceptability and Ad Consistency tasks.
Case Study on the Model’s Behavior on Ad Similarity Task. We used integrated gradients [18] to visualize the attention of Tohoku BERT, aiming for a deeper understanding of the model’s behavior, which can potentially enhance its performance. Figure 5 shows examples of low and high errors between predictions and the ground truth in the Ad Similarity task. In the first example, the model fails to pay sufficient attention to “online seminar,” which has a similar meaning to “webinar.” However, high attention to the date expression “9/9 (Thursday) at 13:00” allows the model to predict a score close to the correct answer. However, in the second example, the model did not pay enough attention to “English conversation coaching,” which has a similar meaning to “gaining English skills,” resulting in an incorrect prediction of dissimilarity. As we found many such cases apart from the ones mentioned, we believe that the model tends to prioritize surface-level information, especially for entities such as dates, times, and numbers. This may prevent the model from correctly identifying cases that are semantically similar even when the surface-level information differs, and vice versa.
Another instance where we observed shortcomings in the models is in reasoning. For example, the models mistakenly interpret that phrases such as “half price” and “buy 2 get 1 free” have the same meaning, whereas humans easily comprehend. Additionally, pairs of phrases like “Available 24 hours a day, 7 days a week” and “Closed every Wednesday,” and “Property located near the station” and “20-minute walk from the station” often confuse the models in understanding. These types of paraphrasing are commonly used in the ad creation process; however, it is challenging to automatically capture them in a natural dataset. Therefore, a specialized dataset focusing on these phenomena is necessary to enable models to capture them accurately.
6 Related Work
Mita et al. [10] proposed an ad text generation benchmark that aims to evaluate NLG models (Step 2 in Figure 1) with surface- and semantic-level overlap metrics, including BLEU [19], ROUGE [20], and BERTScore [21], along with simple heuristics, including keyword insertion rate [22] and text length. However, this approach is inadequate for real-world applications because: (1) it requires reference texts to evaluate generated texts, making it difficult to evaluate new ads that have no references, and (2) simple heuristics cannot guarantee the quality of the ad text for delivery, as more specialized criteria are required in the actual production, such as appropriate wording, effective appeals, consistency between ad text and product information, and high predicted performance.
CTR is widely used to predict ad performance and make marketing strategies in advance. The CTR prediction task was first studied by Robinson et al. [23] and has been actively researched by Rosales et al. [24], McMahan et al. [25], Chapelle et al. [26], Yan et al. [27] and Kumar et al. [28] with traditional machine learning techniques such as logistic regression-based and factorization machines (FM)-based models [29] in past decades [14]. In recent years, neural network (NN)-based models have performed successfully, such as the CNN-based model [30] by Niu and Hou [13], LSTM-based model [31] by Gharibshah et al. [12], and the combination of FM and NN model by Guo et al. [32]. However, while there are many methods for CTR prediction, there are few studies that evaluate the quality of ad texts at a more granular level. Coarse-level feedback may be insufficient to improve ad quality, as it is unclear where edits or what changes should be made. AdTEC, on the other hand, provides more detailed evaluation results, such as acceptability and consistency, making it easier to integrate into actual AdOps workflows.
7 Conclusion
We defined five tasks to verify the quality of ad texts and built the AdTEC, a large, versatile, and comprehensive benchmark of advertising data constructed for the NLP community, based on real-world AdOps workflows. We conducted evaluations with both PLMs and human evaluators on AdTEC to explore its characteristics, offering insights into practical workflow applications and potential areas for future improvement and research. Our findings suggest that sampling directly from real data generally benefits the model, highlighting the importance for tasks centered on natural language inference and semantic understanding. We hope that the combination of our defined tasks and datasets in this paper will advance research in ad text evaluation, bridging the fields of advertising and NLP, and paving the way for new discoveries and applications.
Appendix A Dataset Documentation
We adhere to the existing dataset documentation framework proposed by Gebru et al. [33] to provide comprehensive information about our dataset.
A.1 Motivation
For what purpose was the dataset created?
This dataset is designed to evaluate the quality of ad texts in multiple aspects from the perspective of practical advertising operations.
Who created the dataset (for example, which team, research group) and on behalf of which entity (for example, company, institution, organization)?
The authors of this paper created the dataset and conducted the research as part of a joint program between CyberAgent and the Nara Institute of Science and Technology for CyberAgent.
Who funded the creation of the dataset?
This research was conducted in the joint research between CyberAgent and the Nara Institute of Science and Technology.
A.2 Composition
What do the instances that comprise the dataset represent (for example, documents, photos, people, countries)?
The dataset represents documents, e.g., ad texts, LP texts, keywords, and industry types.
How many instances are there in total (of each type, if appropriate)?
Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set?
As described in 3.2, the dataset includes only a subset of the original data, sampled over a specific period and refined through preprocessing.
What data does each instance consist of?
Our dataset primarily comprises ad text data and related information, including LP texts, keywords, and industry types.
Is there a label or target associated with each instance?
Yes. See Section 3.
Is any information missing from individual instances?
N/A.
Are relationships between individual instances made explicit (for example, users’ movie ratings, social network links)?
N/A.
Are there recommended data splits (for example, training, development/validation, testing)?
Are there any errors, sources of noise, or redundancies in the dataset?
The use of automatically generated ad texts by NLG models may introduce unnatural expressions, which can be considered a form of noise.
Is the dataset self-contained, or does it link to or otherwise rely on external resources (for example, websites, tweets, other datasets)?
The dataset is self-contained for the tasks described in the paper.
Does the dataset contain data that might be considered confidential (for example, data that is protected by legal privilege or by doctor–patient confidentiality, data that includes the content of individuals’ non-public communications)?
Data originally intended for internal use but subsequently made public through appropriate channels.
Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety?
No.
Does the dataset identify any subpopulations (for example, by age, gender)?
No.
Is it possible to identify individuals (that is, one or more natural persons), either directly or indirectly (that is, in combination with other data) from the dataset?
No.
Does the dataset contain data that might be considered sensitive in any way (for example, data that reveals race or ethnic origins, sexual orientations, religious beliefs, political opinions or union memberships, or locations; financial or health data; biometric or genetic data; forms of government identification, such as social security numbers; criminal history)?
No.
A.3 Collection Process
How was the data associated with each instance acquired? Was the data directly observable (for example, raw text, movie ratings), reported by subjects (for example, survey responses), or indirectly inferred/ derived from other data (for example, part-of-speech tags, model-based guesses for age or language)?
See Section 3.2.
What mechanisms or procedures were used to collect the data (for example, hardware apparatuses or sensors, manual human curation, software programs, software APIs)?
See Section 3.2.
If the dataset is a sample from a larger set, what was the sampling strategy (for example, deterministic, probabilistic with specific sampling probabilities)?
As described in Section 3.2, we utilized data from a specific period without employing probabilistic sampling. Items that could not be technically included in the dataset were removed.
Who was involved in the data collection process (for example, students, crowdworkers, contractors) and how were they compensated (for example, how much were crowdworkers paid)?
Authors and other full-time employees. All workers involved in this research are employed as full-time employees by the company and are compensated at rates above the minimum wage set by local authorities. Additionally, they are provided with basic social security benefits.
Over what timeframe was the data collected?
The data collection was conducted in 2023.
Were any ethical review processes conducted (for example, by an institutional review board)?
No.
Did you collect the data from the individuals in question directly, or obtain it via third parties or other sources (for example, websites)?
In question directly.
Were the individuals in question notified about the data collection?
Yes.
Did the individuals in question consent to the collection and use of their data?
Yes.
If consent was obtained, were the consenting individuals provided with a mechanism to revoke their consent in the future or for certain uses?
N/A.
Has an analysis of the potential impact of the dataset and its use on data subjects (for example, a data protection impact analysis) been conducted?
N/A.
A.4 Preprocessing/cleaning/labeling
Was any preprocessing/cleaning/labeling of the data done (for example, discretization or bucketing, tokenization, part-of-speech tagging, SIFT feature extraction, removal of instances, processing of missing values)?
Yes. See Section 3.2.
Was the “raw” data saved in addition to the preprocessed/cleaned/ labeled data (for example, to support unanticipated future uses)?
Yes.
Is the software that was used to preprocess/clean/label the data available?
No.
A.5 Uses
Has the dataset been used for any tasks already?
Yes. The dataset has been used to evaluate the quality of ad texts at CyberAgent.
Is there a repository that links to any or all papers or systems that use the dataset?
See Appendix C.1.
What (other) tasks could the dataset be used for?
Various tasks using the dataset can be considered, such as language modeling on the acceptable ad texts and hallucination detection with the Ad Consistency dataset.
Is there anything about the composition of the dataset or the way it was collected and preprocessed/cleaned/labeled that might impact future uses?
As mentioned in Section 4, the Ad Performance Estimation task masks certain entities to protect information. Consequently, users may need to complete the task with incomplete information, which could affect its usability.
Are there tasks for which the dataset should not be used?
N/A.
A.6 Distribution
Will the dataset be distributed to third parties outside of the entity (for example, company, institution, organization) on behalf of which the dataset was created?
Yes.
How will the dataset be distributed (for example, tarball on website, API, GitHub)?
See Appendix C.1.
When will the dataset be distributed?
After this paper is accepted, as soon as possible.
Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)?
The dataset is distributed under Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International (CC BY-NC-SA 4.0) License.
Have any third parties imposed IP-based or other restrictions on the data associated with the instances?
The third party data is not intended for commercial use and is subject to the organization’s terms and conditions.
Do any export controls or other regulatory restrictions apply to the dataset or to individual instances?
No
A.7 Maintenance
Who will be supporting/hosting/maintaining the dataset?
The authors will be.
How can the owner/curator/ manager of the dataset be contacted (for example, email address)?
By the email address.
Is there an erratum?
N/A.
Will the dataset be updated (for example, to correct labeling errors, add new instances, delete instances)?
It is possible to update the dataset on the website or codebase to correct errors or delete instances upon request.
If the dataset relates to people, are there applicable limits on the retention of the data associated with the instances (for example, were the individuals in question told that their data would be retained for a fixed period of time and then deleted)?
N/A.
Will older versions of the dataset continue to be supported/hosted/ maintained?
It depends on the nature of the dataset update. We may inform dataset users through our website or codebase.
If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so?
No. We consider that the dataset should be used exclusively for our task. Any further work should be a separate contribution from ours.
Appendix B Limitations
Language
The language is limited to Japanese as the advertising agency company is located in Japan. Expressions related to online advertising vary according to language and culture. In Japan, for example, there is a cultural preference for cluttered designs [34], resulting in a lot of information being scattered around compared to the clean LPs in the US and Europe. The search queries are also different: people ask “Which car insurance is the cheapest?” in English-speaking countries, whereas short phrases, such as “cheap car insurance,” are commonly used in Japan. This indicates that different strategies and models are required for different languages and countries, and therefore, benchmarks based on different languages need to be established.
Appendix C Additional Information in Dataset Construction
C.1 Code and Data Availability
The code and data are available at https://github.com/CyberAgentAILab/AdTEC.
C.2 Annotation Guidelines
C.3 Label Distributions
The label distributions for Ad Acceptability, Ad Consistency, A3 Recognition, Ad Similarity, and Ad Performance Estimation tasks are shown in Table 13 and Figure 9.
| Label | Train | Dev | Test | Total |
| Ad Acceptability | ||||
| Acceptable | 15,099 | – | 850 | 15,949 |
| Not acceptable | 20,278 | – | 1,150 | 21,428 |
| Total | 35,377 | – | 2,000 | 37,377 |
| Ad Consistency | ||||
| Consistent | 8,708 | – | 620 | 9,328 |
| Not consistent | 20,048 | – | 1,380 | 21,428 |
| Total | 28,756 | – | 2,000 | 30,756 |
| A3 Recognition | ||||
| Special deals | 230 | 63 | 42 | 335 |
| Discount price | 80 | 22 | 16 | 118 |
| Reward points | 50 | 17 | 14 | 81 |
| Free | 283 | 66 | 62 | 411 |
| Special gift | 83 | 20 | 18 | 121 |
| Features | 746 | 178 | 190 | 1,114 |
| Quality | 35 | 17 | 11 | 63 |
| Problem solving | 10 | 4 | 3 | 17 |
| Speed | 95 | 23 | 20 | 138 |
| User-friendliness | 213 | 65 | 32 | 310 |
| Transportation | 53 | 17 | 10 | 80 |
| Limited offers | 36 | 10 | 6 | 52 |
| Limited time | 43 | 12 | 5 | 60 |
| Limited target | 81 | 10 | 18 | 109 |
| First-time limited | 16 | 5 | 4 | 25 |
| Performance | 47 | 13 | 9 | 69 |
| Largest/No.1 | 108 | 11 | 20 | 139 |
| Product lineup | 167 | 38 | 42 | 247 |
| Trend | 69 | 15 | 13 | 97 |
| Others | 117 | 33 | 23 | 173 |
| Story | 73 | 16 | 9 | 98 |
| Total | 2,635 | 655 | 567 | 3,857 |
| A3 Recognition (#Labels per document) | ||||
| 0 | 337 | 94 | 84 | 515 |
| 1 | 769 | 198 | 165 | 1,132 |
| 2 | 485 | 98 | 100 | 683 |
| 3 | 182 | 44 | 47 | 273 |
| 4 | 69 | 26 | 10 | 105 |
| 5 | 10 | 5 | 3 | 18 |
| 6 | 4 | 0 | 1 | 5 |
| Total | 1,856 | 465 | 410 | 2,731 |
| Ad Similarity | ||||
| 527 | 66 | 67 | 660 | |
| 845 | 105 | 108 | 1,058 | |
| 2,739 | 343 | 344 | 3,426 | |
| 790 | 99 | 100 | 989 | |
| 79 | 10 | 10 | 99 | |
| Total | 4,980 | 623 | 629 | 6,232 |
Appendix D Additional Information in Experiment
D.1 Model Details
The versions and URLs of the models we used are listed in Table 8. The detailed settings for the pretraining of fine-tuned models, such as the tokenizer and dataset, are provided in Table 9.
| Name | URL |
| Tohoku BERT | tohoku-nlp/bert-{base,large}-japanese-v2 |
| Waseda RoBERTa | nlp-waseda/roberta-{base,large}-japanese-with-auto-jumanpp |
| XLM-RoBERTa | xlm-roberta-{base,large} |
| CALM2 | cyberagent/calm2-7b-chat |
| ELYZA | elyza/ELYZA-japanese-CodeLlama-7b-instruct |
| OpenAI | 2023-08-01-preview |
| Model | PreTokenizer (Dictionary) | Tokenization Unit | Dataset | #Vocab | #Param |
| Tohoku BERTBASE | MeCab (IPADic+NEologd) | BPE | Wikipedia (Ja) | 32K | 111M |
| Tohoku BERTLARGE | MeCab (IPADic+NEologd) | BPE | Wikipedia (Ja) | 32K | 337M |
| Waseda RoBERTaBASE | Juman++ | Unigram LM | Wikipedia (Ja) + CC (Ja) | 32K | 110M |
| Waseda RoBERTaLARGE | Juman++ | Unigram LM | Wikipedia (Ja) + CC (Ja) | 32K | 336M |
| XLM-RoBERTaBASE | — | Unigram LM | Multilingual CC | 250K | 278M |
| XLM-RoBERTaLARGE | — | Unigram LM | Multilingual CC | 250K | 559M |
D.2 Hyperparameters
We used PyTorch [35] and HuggingFace [36] libraries for the model implementation. We present the hyperparameters employed during the training of both fine-tuned encoder models and zero-/few-shot settings on LLMs. Table 12 details the specific hyperparameters used in our experiments.
| Parameter | Value |
| Fine-tuned Model | |
| Lerning Rate | {2e-5, 5.5e-5, 2e-6} |
| seed | 0 |
| Epoch | 30 |
| Early Stopping Patience | 3 |
| Optimizer | Adam |
| Max Sequence Length | 128 |
| Large Language Model | |
| Attempts Per an Instance | 5 |
| Temperature | 0.8 |
| Max New Tokens | 64 |
D.3 Prompts
The template for the prompt in each task is illustrated in Figure 7. We begin by outlining the purpose and overview of the task, followed by an explanation of the criteria and standards for evaluation. Next, we provide instructions on the response format, and in some cases, additional few-shot examples are included. Tables 14 and 15 show the prompts actually used in the tasks. Terms in curly brackets are variables to be filled in. All variables, except {few_shot_examples}, correspond to the fields described in Section 3. Note that all prompts are translated from Japanese into English for the presentation brevity.
| Ad Acceptability |
| Based on the following settings, determine whether the input sentence is acceptable as an advertisement expression. |
| # Criteria |
| - Output acceptable if the sentence is coherent, fluent, and easy to read. |
| - Output unacceptable if the sentence is unnatural due to unnecessary and excessive repetition, |
| inappropriate use of symbols, etc. |
| # Formats |
| - Answer with either acceptable or unacceptable. |
| - Do not output anything other than your answer. |
| {few_shot_examples} |
| Input sentence: {ad_text} |
| Output: |
| Ad Consistency |
| Based on the following settings, determine whether the given input sentence is consistent with the LP text. |
| # Criteria |
| - Output inconsistent if the input sentence contains expressions not included in the LP text. |
| - Output inconsistent if the input sentence outputs different numbers or names from those mentioned in the LP text. |
| # Formats |
| - Answer with either consistent or inconsistent. |
| - Do not output anything other than your answer. |
| {few_shot_examples} |
| Input sentence: {ad_text} |
| LP text: {lp_text} |
| Output: |
| Ad Performance Estimation |
| Based on the following settings, estimate the performance of the given advertisement information. |
| # Criteria |
| - The good quality of the ad text (whether it is attractive, whether it is effective in the industry) |
| - The relevance of keywords and ad text |
| - The relevance of LP text and ad text |
| # Formats |
| - Do not output anything other than your answer. |
| - Answer with a number between 0 and 100. |
| {few_shot_examples} |
| Industry: {industry_type} |
| Keyword: {keyword} |
| Title: {title_1} {title_2} {title_3} |
| Description: {description_1} {description_2} |
| Output: |
| Ad Similarity |
| Based on the following settings, rate the advertising similarity of the given two input sentences. |
| # Criteria |
| - Judge based on the similarity in product category/product name and the similarity in appeal axis/persuasive |
| expressions. |
| - Closer to 5 if similar, closer to 1 if not similar. |
| # Formats |
| - Do not output anything other than your answer. |
| - Answer with a decimal number between 1 and 5. |
| {few_shot_examples} |
| Input sentence 1: {ad_text_1} |
| Input sentence 2: {ad_text_2} |
| Output: |
UTF8ipxm
| A3 Recognition |
| Based on the following settings, list all the aspects of advertising appeals included in the given input sentence. |
| Choose aspect labels from the following list. The list is in the format of "Label": "Description". |
| # List of aspects of advertising appeals |
| - Special deals: Expressions emphasizing a sense of value such as price or discounts |
| - Discount price: Expressions emphasizing specific prices or discounts |
| - Reward points: Expressions emphasizing the rebate of points or money |
| - Free: Expressions emphasizing that something is free |
| - Special gift: Expressions emphasizing that perks are included |
| - Features: Expressions emphasizing the content or features of the service |
| - Quality: Expressions emphasizing high quality or a high-grade feel |
| - Problem solving: Expressions emphasizing the solution to customers’ problems |
| - Speed: Expressions emphasizing the speed of delivery or procedures |
| - User-friendliness: Expressions emphasizing the ease of use of the service |
| - Transportation: Expressions emphasizing good accessibility |
| - Limited offers: Expressions emphasizing some form of limitation |
| - Limited time: Expressions emphasizing a limited period for service provision |
| - Limited target: Expressions emphasizing that the service is provided to/for a limited target |
| - First-time limited: Expressions emphasizing that the service is provided only for the first time |
| - Performance: Expressions emphasizing the achievements of the service or company |
| - Largest/No.1: Expressions emphasizing the scale or being No. 1 of the service or company |
| - Product lineup: Expressions emphasizing the assortment or the number of stores |
| - Trend: Expressions emphasizing that it is trending or in vogue |
| - Other: Persuasive expressions suitable for advertising that do not fall into the above labels |
| - Story: Expressions emphasizing the synopsis of the work |
| - No Match: Label for when there is no persuasive expression |
| # Formats |
| - Output each aspects separated by “|” |
| - Do not output anything other than your answer |
| - If no aspects apply, output “No Match” |
| {few_shot_examples} |
| Input sentence: {ad_text} |
| Output: |
D.4 Complete Version of Table 6
| Evaluator | Ad Accept. | Ad Consist. | Ad Perform. Est. | A3 Recognition | Ad Similarity |
| Accuracy/F1-score | Accuracy/F1-score | Pearson/Spearman | F1-micro/-macro | Pearson/Spearman | |
| Fine-tuned Encoder Models | |||||
| Tohoku BERTBASE | 0.685/0.691 | 0.757/0.504 | 0.437/0.454 | 0.753/0.629 | 0.769/0.803 |
| Tohoku BERTLARGE | 0.711/0.688 | 0.767/0.552 | 0.480/0.497 | 0.774/0.694 | 0.773/0.807 |
| Waseda BERTBASE | 0.615/0.639 | 0.725/0.388 | 0.444/0.454 | 0.641/0.442 | 0.749/0.797 |
| Waseda BERTLARGE | 0.598/0.637 | 0.755/0.474 | 0.445/0.457 | 0.663/0.517 | 0.740/0.800 |
| XLM-RoBERTaBASE | 0.694/0.677 | 0.743/0.465 | 0.425/0.439 | 0.730/0.542 | 0.846/0.870 |
| XLM-RoBERTaLARGE | 0.705/0.690 | 0.758/0.519 | 0.453/0.457 | 0.778/0.677 | 0.878/0.878 |
| Zero-/Few-shot LLMs | |||||
| CALM27b | 0.520/0.115 | 0.381/0.472 | 0.006/0.013 | 0.154/0.042 | 0.036/0.036 |
| ELYZA7b | 0.352/0.520 | 0.628/0.771 | 0.003/0.046 | 0.196/0.044 | 0.015/-0.004 |
| GPT-3.5 | 0.369/0.489 | 0.528/0.570 | -0.013/-0.022 | 0.255/0.064 | 0.389/0.385 |
| GPT-4 | 0.325/0.433 | 0.583/0.612 | 0.028/0.073 | 0.417/0.113 | 0.776/0.811 |
| Human | 0.732/0.790 | 0.703/0.807 | — | 0.564/0.538 | 0.699/0.765 |